Verify that the given family of functions solves the differential equation. (i)
dy dt
(ii)
dy dt
= (1 − 2t)y2, y = (c) Evaluate the integral
= y2 sin t, y = 1
0 ( 1
C − t + t2 . 1
C + cos t x2
√ . 4 − x2)3 dx
1
Expert's answer
2020-06-29T17:12:44-0400

(i) Given "\frac{dy}{dt} = (1-2t)y^2" and "y = \frac{1}{c-t+t^2}"

Differentiating y with respect to x,

"\frac{dy}{dx} = -(\frac{1}{c-t+t^2})^2(-1+2t) = (\frac{1}{c-t+t^2})^2(1-2t)"

replacing value of y

we obtain, "\frac{dy}{dt} = (1-2t)y^2"


(ii) Given "\frac{dy}{dt} = y^2sin(t)" and "y = \frac{1}{c+cost}"

differentiating both sides with respect to x,

"\frac{dy}{dt} = -(\frac{1}{c+cost})^2(-sint) = (\frac{1}{c+cost})^2(sint)"

replacing value of y with y

we obtain, "\frac{dy}{dt} = y^2sin(t)"


(iii) We need to find integration of "\int_0^1 x^2 (\sqrt{4-x^2})^3dx"

let "x = 2sin(u)"

then "dx = 2cos(u) du"

Let us solve integration first without limit

"\int 8cos(u)sin^2(u)(4-4sin^2(u))^{3/2}du"


"=\int 8cos(u)sin^2(u)(4cos^2(u))^{3/2}du"

"= \int 64cos^4(u)sin^2(u)du"

"= \int 64cos^6(u)(1-cos^2(u))du"

integrating it we get,

"-\frac{x(4-x^2)^{5/2}}{6} + \frac{x(4-x^2)^{3/2}}{6}+x(4-x^2)^{1/2} + 4asin(\frac{x}{2}) + C"

putting the limits, we get

"\int_0^1 x^2 (\sqrt{4-x^2})^3dx = \frac{2\pi}{3}"